# UMA: Ultra-detailed Human Avatars via Multi-level Surface Alignment

### v1

@@ 3.5 Texel Level Alignment
Through the depth and vertex-level alignment, we can now generate a drivable template mesh $M(\bm{\bar{\theta}}_{f},\mathbf{z}_{f})=\mathbf{V}_{f}$, which is closely attached to the ground truth surface $\mathbf{V}_{\mathrm{gt},f}$ and in correspondence over time. Conditioned on this template $\mathbf{V}_{f}$, we can further train the detail dynamic human appearance represented with animatable Gaussian textures $\mathbf{T}^{\mathrm{gs}}_{f}$ by minimizing $\mathcal{L}_{\mathrm{gau}}$ (Eq. 5). However, the Gaussian texels may also drift on the surface hindering learning the highest frequency detail (see the discussion in Sec. 3.2). Thus, we introduce an additional texel level correspondence loss
$\mathcal{L}_{\mathrm{corr-tex}}=\sum_{f=1}^{F}\sum_{i\in\mathcal{V}^{\mathrm{tex}}_{f}}||\bm{\mu}_{i}-\tilde{\mathbf{P}}^{\mathrm{tex}}_{f,i}||_{2}^{2},$ (12)
which supervises the Gaussian positions based on their correspondence $\tilde{\mathbf{P}}^{\mathrm{tex}}_{f,i}$. The derivation of $\tilde{\mathbf{P}}^{\mathrm{tex}}_{f,i}$ is analogous to the one for vertices (Sec. 3.4) simply adding barycentric interpolation. A detailed explanation is provided in the supplemental material. Thus, we can even compensate for finest level drifts further improving detail preservation (see Sec. 4.2).
4 Results
@@
Implementation Details. Our approach is implemented with PyTorch [68]. For rendering, we adopt Analytical Splatting [64] for rasterizing the 3D Gaussian splats and employ Nvdiffrast [67] to render the textured meshes. All the stages, competing methods and ablation alternatives are trained and tested on a server with two NVIDIA H100 graphics cards and a AMD EPYC 9554 CPU. The depth alignment stage is trained for $360{,}000$ iterations using the Adam optimizer [69] with a learning rate of $41e-$ 4 scheduled by a cosine decay, which takes around 12 hours. The vertex alignment stage is trained for $360{,}000$ iterations using Adam optimizer with a learning rate of $45e-$ 4 scheduled with a cosine decay scheduler, which takes around 12 hours. The resolution of the Gaussian textures was set to $768\times 768$ for all the subjects, resulting in roughly $250k$ Gaussian splats in total. To train the animatable Gaussian texture with texel-level alignment, following the open-sourced implementation in Pang et al.[7], it includes $15{,}000$ iterations of initialization before the main training, while the main training lasts for $2{,}000{,}000$ iterations with a learning rate of $1e^{-4}$. The model is trained at an image resolution of $1620\times 3072$ on random crops of size $810\times 1536$.
@@
Dataset. Our new dataset features five subjects wearing a diverse range of apparel, including loose-fitting tops and skirts. Notably, unlike previous datasets that often include garments with limited texture complexity or solid colors, the subjects in our dataset wear clothing with rich and intricate patterns. For each subject, we captured separate training and testing sequences in which they perform a variety of everyday motions, including jumping jacks, dancing, and boxing. The sequences are captured using a multi-camera system consisting of $40$ synchronized and calibrated cameras, each recording at a resolution of $3240\times 6144$ and a frame rate of $25$ fps. The training sequences comprise approximately $17{,}000$ frames, while the testing sequences contain around $7{,}000$ frames. Each frame of the captured videos is annotated with skeletal poses obtained using commercial 3D pose estimation software [60], foreground segmentation generated by Sapiens [70], and pseudo ground-truth geometry reconstructed with NeuS2 [61]. Moreover, to facilitate comparison with existing methods, we additionally provide SMPL-X [42] parameters for the video frames.
@@
Metrics. We adopt the Peak Signal-to-Noise Ratio (PSNR) metric to measure the quality of the rendered image. Besides, we adopt the Structural Similarity Index (SSIM) and learned perceptual image patch similarity (LPIPS)  [71] that better mirrors human perception. Note that the metrics are computed at a resolution of $1620\times 3072$, averaged over every 10th frame in the testing sequences, using two camera views that were excluded during training. To access the geometry reconstruction accuracy, we compute the Chamfer distance between generated mesh vertices (or 3D Gaussian point clouds) and the pseudo ground-truth reconstructions.
Qualitative Results. Fig. 4 demonstrates diverse qualitative rendering results of our method on different subjects and diverse motions. Our method faithfully recovers finest cloth wrinkles and cloth patterns for, both, novel view synthesis and novel pose synthesis tasks.
@@
Figure 4: Qualitative Results. UMA performs well on, both, novel view and novel pose synthesis tasks, and captures extremely high-resolution details on human avatars, e.g., texture patterns and cloth wrinkles.
4.1 Comparisons
@@
DDC. [72] 30.54 0.9215 112.3 1.579 26.94 0.8965 127.6 1.817
MeshAvatar [39] 27.23 0.8873 105.7 1.898 25.98 0.8805 117.1 2.393
TriHuman [49] 30.24 0.9166 83.04 0.983 27.37 0.8977 106.3 1.495
ASH [7] 35.96 0.9569 63.84 2.100 27.50 0.8974 112.4 2.377
Animatable Gaussians [59] 29.07 0.9042 75.42 2.323 26.06 0.8839 103.3 2.923
GaussianAvatar [58] 25.88 0.8884 127.2 3.273 25.26 0.8845 134.9 3.630
3DGS-Avatar [55] 25.55 0.8865 141.6 3.876 24.87 0.8822 146.7 4.129
Ours 36.78 0.9655 42.14 0.899 27.64 0.8943 90.60 1.517
TABLE I: Quantitative Comparison. Here, we quantitatively compare our method with prior works on rendering quality and geometry accuracy on the training split and testing split across all sequences. We highlight the best, second-best, and third-best scores. UMA consistently outperforms previous methods in all metrics, especially for the perceptual matrix which better reflect the reconstruction of fine details.
@@
Mesh-based Approaches. MeshAvatar [39] models the pose-dependent surface deformations and materials conditioned from orthogonal projected position maps. DDC [72] models motion-aware clothing dynamics using learned deformation and pose-conditioned texture maps.
Implicit-based Approaches. TriHuman [49] is an implicit-based method that models dynamic geometry and appearance of the dynamic clothed human via a signed distance field and a color field, both conditioned on motion-dependent tri-planes in tangent space. It employs unbiased volume rendering [73] to couple geometry and appearance fields.
Point-based Approaches. ASH [7], 3DGS-Avatar [55], and GaussianAvatar [58] model detailed appearance using pose-dependent 3D Gaussians Splats defined in the texel space. 3DGS-Avatar [55] and GaussianAvatar [58] employ a parametric human body model [42] to represent coarse geometry, while ASH [7] leverages learned embedded deformations of the person-specific template mesh inspired by DDC [72]. Notebly, we set the resolution of the Gaussian textures of ASH to $768$ for fair comparison. Animatable Gaussians [59] models coarse-level geometry using a person-specific template mesh and represents appearance using 3D Gaussian splats inferred from a front-and-back orthogonally projected position map.
@@
Ours 38.28 0.9689 27.09 0.815 29.66 0.9076 65.56 1.112
TABLE II: Ablation Study. We compare our full method with design alternatives. We highlight the best, second-best, and third-best scores. By addressing depth misalignment as well as surface drift, and we observe consistent improvements in view synthesis and geometry generation accuracy across, both, the training and testing splits of the dataset.
@@
Quantitative Comparison. Tab. I presents the quantitative comparison against competing methods for novel view and novel pose synthesis. For novel view synthesis, our method consistently outperforms competing methods in all metrics, with a particularly significant improvement in LPIPS, which better reflects human perception. For novel pose synthesis, UMA achieves the highest PSNR and LPIPS scores, highlighting its robustness to unseen poses. For geometry synthesis, UMA attains the lowest Chamfer distance on the training split and ranks second on the testing split with only a marginal difference. Thanks to our multi-level alignments, UMA achieves significant improvements over all Gaussian-based methods.
@@
Qualitative Comparison. Fig. 5 presents the qualitative comparison on novel-view and pose renderings. DDC [10] and TriHuman [49] can not recover sharp texture patterns limited by their appearance representations. ASH [7] and Animatable Gaussians [59] could recover high frequency details. However, due to the depth misalignment and surface drift, they tend to produce blurry and wrong texture patterns. In stark contrast, UMA is able to reproduce more accurate and sharper texture patterns and cloth wrinkles.
@@
Figure 5: Qualitative Rendering Comparison. We compare our approach with the competing approaches on novel view synthesis and novel pose generation. Compared with other methods, our approach preservers the best levels of details. Please zoom-in to better observe the details.
@@
Figure 6: Qualitative Geometry Comparison. We compare our approach with the competing approaches on novel view synthesis and novel pose generation. Compared with other methods, our approach preservers the best level of detail. Please zoom-in to better observe the details.
4.2 Ablations
@@
To assess the impact of our core design choices, accounting for depth, vertex, and texel alignment, we conduct ablation studies by progressively building upon a baseline model with major components.
@@
Baseline. We begin with the baseline model (raw), namely the animatable Gaussian textures introduced in Sec. 3.1. As shown in Fig. 8, though the baseline model captures pose-dependent clothing wrinkles, it fails to reproduce fine texture patterns on the dress as the depth misalignment caused by the one-to-many mapping between the skeletal motion and surface deformations.
@@
Depth Alignment. To address the depth misalignment, as mentioned in Sec. 3.3, we applied a per-frame latent $\bm{z}_{f}$ as the additional input apart from the skeletal motion $\bm{\bar{\theta}}_{f}$ for predicting the drivable template mesh, termed as raw + lat. As shown in Tab. II, this improvement in depth alignment significantly enhances the accuracy of surface reconstruction, resulting in improved performance on, both, novel view and novel pose synthesis. However, the detailed texture pattern is missing due to the on-surface drift.
@@
Figure 7: Qualitative Ablation on Geometry. Naive Gaussian textures (raw) (A) exhibits highest error due to the depth misalignment. Depth alignment (B) improves the surface accuracy. Comparing to naive correspondence tracking (longt.) (C) and differentiable rendering (rdr.) (D), our vertex alignment (trk. 1st/2st) (E, F) achieves higher surface quality. Our texel level alignment (G) further improves high-frequency surface details.
@@
Figure 8: Qualitative Ablation on Renderings. Our depth alignment (raw + lat) greatly improves the rendering sharpness of wrinkles and texture patterns. Comparing to naive variants (raw + lat + longt. and raw + lat + rdr.), our triangle level alignment (raw + lat + trk. 1st and raw + lat + trk. 2nd) further refines rendering quality. Note the cross patterns in the yellow box. Our full model with texel-level alignment even recovers smallest patterns, i.e., the stripes in the green box.
Vertex-level Alignment. We compare two alternatives for supervising surface correspondence of the template mesh: using differentiable rendering losses (raw + lat + rdr.) and using correspondences extracted from the video point tracker [11] across the entire sequence (raw + lat + logt.). As shown in Fig. 8, directly supervising with tracked correspondences (raw + lat + logt.) leads to blurry results due to accumulated tracking errors over time. Differentiable rendering (raw + lat + rdr.) also offers limited improvement since it is often stuck in local minima. In contrast, our method leverages the animatable template mesh to compute and refine per-frame correspondences (raw + lat + trk. 1st), providing more accurate and temporally consistent supervisions. By leveraging the template mesh supervised with more coherent correspondences, our method preserves finer texture details and significantly improves quantitative performance. Iterating the vertex alignment for a second time (raw + lat + trk. 2st) further improves the results.
@@
Texel-level Alignment. As discussed in Sec. 3.5, texel-level alignment provides dense and precise supervision by establishing detailed correspondences between texels and the ground-truth surface. By leveraging the texel-level correspondences, our full approach (Ours) achieves a substantial improvement in surface reconstruction accuracy, along with further gains in, both, accuracy and visual quality for novel view and pose synthesis.
5 Conclusion and Limitations

### v2

@@ 3.5. Texel-level Alignment
Through the depth and vertex-level alignment, we can now generate a drivable template mesh $M(\boldsymbol{\bar{\theta}}_{f},\mathbf{z}_{f})=\mathbf{V}_{f}$, which is closely attached to the ground truth surface $\mathbf{V}_{\mathrm{gt},f}$ and in correspondence over time. Conditioned on this template $\mathbf{V}_{f}$, we can further train the detail dynamic human appearance represented with animatable Gaussian textures $\mathbf{T}^{\mathrm{gs}}_{f}$ by minimizing $\mathcal{L}_{\mathrm{gau}}$ (Eq. 10). However, similar to the template mesh vertices, the Gaussian texels may also drift on the surface hindering learning the highest frequency detail (see the discussion in Sec. 3.2). Thus, we introduce an additional texel-level correspondence loss $\mathcal{L}_{\mathrm{cor-tex}}$, which serves as an additional regularization term for the position of the 3D Gaussian Splats $\boldsymbol{\mu}_{i}$:
[…]
where $\tilde{\mathbf{P}}^{\mathrm{tex}}_{f,i}\in\mathbb{R}^{2}$ denotes the corresponding position for the $i$th texel on the ground truth surface. The derivation of $\tilde{\mathbf{P}}^{\mathrm{tex}}_{f,i}$ is analogous to the one for vertices (Sec. 3.4). We can now even compensate for finest level drifts of the Gaussian textures by supervising them on a combined loss
(24) $\mathcal{L}_{\mathrm{gau}}+\mathcal{L}_{\mathrm{cor-tex}},$
further improving detail preservation (see Sec. 4.3).
3.6. Texel Super Resolution
Increasing the resolution of the Gaussian Textures, i.e., the number of the 3D Gaussian Splats, could effectively improve the quality of modeling the fine structures, for example, the yarn patterns on the clothing. However, directly tracking and learning high resolution Gaussian Textures is slow and memory extensive. Therefore, rather than learn the highest resolution directly, we propose a light-weight Gaussian Texture super-resolution module $\mathcal{E}_{\mathrm{sr}}$ to further improve the fidelity of the Gaussian Texture. The Gaussian texture super-resolution module takes the Gaussian Texture $\mathbf{T}^{\mathrm{gs}}_{f}$ from the geometry and appearance network, and produces the up-sampled Gaussian Texture $\mathbf{T}^{\mathrm{gs},*}_{f}$ represented with the residual w.r.t., the original textures:
(25) $\mathbf{T}^{\mathrm{gs},*}_{f}=\mathcal{E}_{\mathrm{sr}}(\mathbf{T}^{\mathrm{gs}}_{f})+f_{\mathrm{int}}(\mathbf{T}^{\mathrm{gs}}_{f})$
where $f_{\mathrm{int}}$ denotes the bilinear interpolation, which produces the Gaussian Texture with doubled resolution. Thanks to the accurate surface tracking over the vertex level and texture level, the light-weight Gaussian texture super-resolution module can focus on the local regions and produces more intricate details such as yarn patterns which can be viewed in a much higher zoomin-level while incurring minimal overhead.
Table 1. Training status for each component. We provide the status for each component and each loss function in the different training stages. The status of each component , indicating that the weights will be updated and the losses are used for training. Depth.A., Vert.A., Tex.A. and Tex.SR. denotes Depth Alignment (Sec. 3.3), Vertex Alignment (Sec. 3.4), Texel Alignment (Sec. 3.5) and Texel Super Resolution (Sec. 3.6), respectively.
Modules Supervision
$\mathcal{F_{\mathrm{eg}}},\mathcal{F_{\mathrm{delta}}}$ $\mathcal{E}_{\mathrm{app}},\mathcal{E}_{\mathrm{geo}}$ $\mathcal{E}_{\mathrm{sr}}$ $\mathcal{L}_{\mathrm{temp}}$ $\mathcal{L}_{\mathrm{gau}}$ $\mathcal{L}_{\mathrm{cor-vrt}}$ $\mathcal{L}_{\mathrm{cor-tex}}$
Depth.A. (Sec. 3.3)
Vert.A. (Sec. 3.4)
Tex.A. (Sec. 3.5)
Tex.SR. (Sec. 3.6)
4. Results
@@
Implementation Details. Our approach is implemented in PyTorch (Paszke et al., 2017). For rendering, we adopt Analytical Splatting (Liang et al., 2024) for rasterizing the 3D Gaussian splats and employ Nvdiffrast (Laine et al., 2020) to render the textured meshes. All the stages, competing methods and ablation alternatives are trained and tested on a server with two NVIDIA H100 graphics cards and a AMD EPYC 9554 CPU. The depth alignment stage is trained for 360,000360{,}000 iterations using the Adam optimizer (Kingma and Ba, 2017) with a learning rate of $1e^{-4}$ scheduled by a cosine decay, which takes around 12 hours. The vertex alignment stage is trained for 360,000360{,}000 iterations using Adam optimizer with a learning rate of $5e^{-4}$ scheduled with a cosine decay scheduler, which takes around 12 hours. The resolution of the Gaussian textures is set to $768\times 768$ for all the subjects, resulting in roughly $250k$ Gaussian splats in total. After the Gaussian super resolution module, the Gaussian textures are upsampled to $1536\times 1536$, i.e., approximately $1$ million Gaussian Splats in total. To train the animatable Gaussian texture with texel-level alignment, following the open-sourced implementation in Pang et al. (2024), it includes 15,00015{,}000 iterations of initialization before the main training, while the main training lasts for 2,000,0002{,}000{,}000 iterations with a learning rate of $1e^{-4}$. The texel super resolution stage takes 1,000,0001{,}000{,}000 iterations with a learning rate of $1e^{-4}$. The model is trained at an image resolution of $1620\times 3072$ on random crops of size $810\times 1536$.
@@
Dataset. Our new dataset features five subjects wearing a diverse range of apparel, including loose-fitting tops and skirts. Notably, unlike previous datasets that often include garments with limited texture complexity or solid colors, the subjects in our dataset wear clothing with rich and intricate patterns. For each subject, we captured separate training and testing sequences in which they perform a variety of everyday motions, including jumping jacks, dancing, and boxing. The sequences are captured using a multi-camera system consisting of $40$ synchronized and calibrated cameras, each recording at a resolution of $3240\times 6144$ and a frame rate of $25$ fps. The training sequences comprise approximately 17,00017{,}000 frames, while the testing sequences contain around 7,0007{,}000 frames. Each frame of the captured videos is annotated with skeletal poses obtained using commercial 3D pose estimation software (TheCaptury, 2020), foreground segmentation generated by Sapiens (Khirodkar et al., 2024), and pseudo ground-truth geometry reconstructed with NeuS2 (Wang et al., 2023). Moreover, to facilitate comparison with existing methods, we additionally provide SMPL-X (Pavlakos et al., 2019) parameters for the video frames.
@@
Metrics. We adopt the Peak Signal-to-Noise Ratio (PSNR) metric to measure the quality of the rendered image. Besides, we adopt the Structural Similarity Index (SSIM) and learned perceptual image patch similarity (LPIPS)  (Zhang et al., 2018) that better mirrors human perception. Note that the metrics are computed at a resolution of $1620\times 3072$, averaged over every 10th frame in the testing sequences, using two camera views that were excluded during training. To assess the geometry reconstruction accuracy, we compute the Chamfer distance between generated mesh vertices (or 3D Gaussian point clouds) and the pseudo ground-truth reconstructions.
4.1. Qualitative Results
Image Synthesis. Fig. 4 presents the image synthesis results. For both tight and loose fitting outfits, UMA faithfully recovers the garment wrinkles finest texture patterns in rendering for both novel views and also under novel poses.
Geometry Synthesis. Additionally, we show the geometry synthesis results in Fig. 5. Note that UMA perseveres the clothing dynamics and the motion-aware detailed deformation of the clothing for the training poses. For the novel motion unseen during training, UMA could produce plausible and vivid clothing dynamics, which is more prominent for loose clothing such as dresses.
@@ 4.1. Qualitative Results
Figure 4. Qualitative Rendering Results. UMA performs well on both novel view and novel pose synthesis tasks, and manages to capture ultra details on human avatars, i.e texture patterns, cloth wrinkles. Please zoom-in to better observe the details.
Figure 5. Qualitative Geometry Results. For both training motions and testing motions unseen during taring, UMA generates clothing with realistic dynamics and vivid detailed deformations. Please zoom-in to better observe the details.
4.2. Comparisons
Table 2. Quantitative Comparison. Here, we quantitatively compare our method with prior works on rendering quality and geometry accuracy on the training split and testing split across all sequences. We highlight the best, second-best, and third-best scores. UMA consistently outperforms previous methods in all metrics, especially for the perceptual matrix which better reflect the reconstruction of fine details.
@@ 4.2. Comparisons
DDC. (Habermann et al., 2021b) 30.54 0.9215 112.3 1.579 26.94 0.8965 127.6 1.817
MeshAvatar (Chen et al., 2024) 27.23 0.8873 105.7 1.898 25.98 0.8805 117.1 2.393
TriHuman (Zhu et al., 2023) 30.24 0.9166 83.04 0.983 27.37 0.8977 106.3 1.495
ASH (Pang et al., 2024) 35.96 0.9569 63.84 2.100 27.50 0.8974 112.4 2.377
Animatable Gaussians (Li et al., 2024) 29.07 0.9042 75.42 2.323 26.06 0.8839 103.3 2.923
GaussianAvatar (Hu et al., 2024) 25.88 0.8884 127.2 3.273 25.26 0.8845 134.9 3.630
3DGS-Avatar (Qian et al., 2024) 25.55 0.8865 141.6 3.876 24.87 0.8822 146.7 4.129
Ours wo SR 36.80 0.9657 41.90 0.876 27.66 0.8943 90.21 1.523
Ours 37.15 0.9681 35.02 0.876 27.68 0.8937 84.12 1.523
@@
Mesh-based Approaches. MeshAvatar (Chen et al., 2024) models the pose-dependent surface deformations and materials conditioned from orthogonal projected position maps. DDC (Habermann et al., 2021b) models motion-aware clothing dynamics using learned deformation and pose-conditioned texture maps.
Implicit-based Approaches. TriHuman (Zhu et al., 2023) is an implicit-based method that models dynamic geometry and appearance of the dynamic clothed human via a signed distance field and a color field, both conditioned on motion-dependent tri-planes in tangent space. It employs unbiased volume rendering (Wang et al., 2021a) to couple geometry and appearance fields.
Point-based Approaches. ASH (Pang et al., 2024), 3DGS-Avatar (Qian et al., 2024), and GaussianAvatar (Hu et al., 2024) model detailed appearance using pose-dependent 3D Gaussians Splats defined in the texel space. 3DGS-Avatar (Qian et al., 2024) and GaussianAvatar (Hu et al., 2024) employ a parametric human body model (Pavlakos et al., 2019) to represent coarse geometry, while ASH (Pang et al., 2024) leverages learned embedded deformations of the person-specific template mesh inspired by DDC (Habermann et al., 2021b). Notably, we set the resolution of the Gaussian textures of ASH to $768\times 768$, which is the same for the texture resolution adopted in our work for fair comparison. Animatable Gaussians (Li et al., 2024) models coarse-level geometry using a person-specific template mesh and represents appearance using 3D Gaussian splats inferred from a front-and-back orthogonally projected position map.
Table 3. Ablation Study. We compare our full method with design alternatives. By addressing depth misalignment as well as surface drift, and we observe consistent improvements in view synthesis and geometry generation accuracy across both the training and testing splits of the dataset.
@@
Ours wo SR 38.38 0.9697 25.87 0.771 29.74 0.9075 63.62 1.151
Ours 38.56 0.9704 25.42 0.771 29.74 0.9076 62.71 1.151
@@
Quantitative Comparison. Tab. 2 presents the quantitative comparison against competing methods for novel view and novel pose synthesis. For novel view synthesis, our method consistently outperforms competing methods in all metrics, with a particularly significant improvement in LPIPS, which better reflects human perception. For novel pose synthesis, UMA achieves the highest PSNR and LPIPS scores, highlighting its robustness to unseen poses. For geometry synthesis, UMA attains the lowest Chamfer distance on the training split and ranks second on the testing split with only a marginal difference to TriHuman. Thanks to our multi-level alignments, UMA achieves significant improvements over all Gaussian-based methods.
@@
Qualitative Comparison. Fig. 6 presents the qualitative comparison on novel view and pose renderings. DDC (Habermann et al., 2021a) and TriHuman (Zhu et al., 2023) can not recover sharp texture patterns limited by their appearance representations. ASH (Pang et al., 2024) and Animatable Gaussians (Li et al., 2024) could recover high frequency details. However, due to the depth misalignment and surface drift, they tend to produce blurry and wrong texture patterns. In stark contrast, UMA is able to reproduce more accurate and sharper texture patterns and cloth wrinkles.
@@
Figure 6. Qualitative Rendering Comparison. We compare our approach with the competing approaches on novel view synthesis and novel pose generation. Compared with other methods, our approach preservers the best levels of details. Please zoom-in to better observe the details. We refer to the supplemental document and video for additional qualitative comparisons with more methods and for the dynamic results.
@@
Figure 7. Qualitative Geometry Comparison. We compare our approach with the competing approaches on novel view synthesis and novel pose generation. Compared with other methods, our approach preservers the best level of detail. Please zoom-in to better observe the details.
4.3. Ablations
@@
To assess the impact of our core design choices, accounting for depth, vertex, and texel alignment, as well as the texel super-resolution, we conduct ablation studies by progressively building upon a baseline model with major components.
@@
Baseline. We begin with the baseline model (raw), namely, the animatable Gaussian textures introduced in Sec. 3.1. Note that we adopt the Analytical Splatting (Liang et al., 2024) for rendering the Gaussaian textures for all the ablative experiments, which has proven capability on rendering at different scales. As shown in Fig. 9, though the baseline model captures pose-dependent clothing wrinkles, it fails to reproduce fine texture patterns on the dress as the depth misalignment caused by the one-to-many mapping between the skeletal motion and surface deformations.
@@
Depth Alignment. To address the depth misalignment, as mentioned in Sec. 3.3, we applied a per-frame latent $\boldsymbol{z}_{f}$ on top of the baseline approach as the additional input apart from the skeletal motion $\boldsymbol{\bar{\theta}}_{f}$ for predicting the drivable template mesh, termed as raw + lat. As shown in Tab. 3, this improvement in depth alignment effectively enhances the accuracy of surface reconstruction. More specifically, by applying the latent conditioning, the one-to-many ambiguity for the coarse template geometry, which attributes to the large scale dynamics error, is addressed. The improved template geometry also attributes to the improved performance on, both, novel view and novel pose synthesis. However, as illustrated in Fig. 9, the absence of explicit surface correspondence supervision leads to on-surface drift, causing fine texture patterns to be missing in both novel-view and novel-pose renderings.
@@
Figure 8. Qualitative Ablation on Geometry. Naive Gaussian textures (raw) (A) exhibits highest error due to depth misalignment. (B) Introducing learnable latent(raw + lat) improves surface reconstruction accuracy. Comparing to naive correspondence tracking (raw + lat + longt.) (C) and differentiable rendering (raw + lat + rdr.) (D), our vertex-level alignment (trk. 1st/2nd) (E, F) achieves much higher accuracy. Our model (Ours wo SR) with texel-level alignment (G) achieves the highest reconstruction fidelity.
@@
Figure 9. Qualitative Ablation on Renderings. Depth alignment (raw + lat) enhances the fidelity of wrinkles and texture patterns. Compared with other alternatives (raw + lat + longt. and raw + lat + rdr.), our approach with vertex-level alignment (raw + lat + trk. 1st and raw + lat + trk. 2nd) further improves rendering quality, see the crossing highlighted in the yellow box. Finally, our model (Ours wo SR) with texel-level alignment recovers even the finest appearance details, such as the dots shown in the green box. Please zoom-in to better observe the details.
Vertex-level Alignment. We compare two alternatives for supervising surface correspondence of the template mesh: using differentiable rendering losses (raw + lat + rdr.) and using correspondences directly extracted from the video point tracker (Karaev et al., 2024) across the entire sequence (raw + lat + logt.).
To supervise surface correspondence with a differentiable rendering loss (raw + lat + rdr.), we augment the drivable template loss $\mathcal{L}_{\mathrm{temp}}$ by rendering the static textures $\mathbf{T}_{\mathrm{0}}$ with Nvdiffrast (Laine et al., 2020) and applying a $L1$ photometric loss between the rendered static-textured template mesh and the ground truth image. As shown in Tab. 3 and Fig. 9, applying differentiable rendering (raw + lat + rdr.) yields only limited quantitative and qualitative improvements, as the optimization often gets trapped in local minima caused by diverse shading and surface deformations.
As another design alternative, we employ tracked correspondences (raw + lat + logt.) obtained directly from a video point tracker from different views as supervision. Specifically, the vertices of the posed template mesh in the first frame are projected into all training camera views and serve as the initialization for point tracking in each view. Then, we adopt the video point tracker to track the point movements starting from the first frame to the end of the sequence. The tracked points/vertices in the image space are unprojected to 3D to served as the supervision for the template mesh vertices. As is shown in Fig. 9 due to the accumulated errors, the naively tracked correspondence could not provide meaningful supervision, and leads to even more blurry results.
In stark contrast, our approach leverages the animatable template mesh to aggregate and refine correspondences (raw + lat + trk. 1st) across multiple views and frames, yielding more accurate correspondence supervision for the template vertices. This more accurate and temporally consistent correspondence supervision is reflected in the recovery of fine texture details that would otherwise appear blurred without proper correspondence supervision, as shown in Fig. 9. Moreover, as shown in Tab. 3, the quantitative performance is significantly improved for both novel-view and novel-pose rendering tasks, with especially large gains in the LPIPS metric, which better reflects human perception.
We further iterate the vertex alignment for a second time (raw + lat + trk. 2st), which further improves both qualitative and quantitative performance.
@@
Texel-level Alignment. As discussed in Sec. 3.5, texel-level alignment provides dense and precise supervision by establishing detailed correspondences between texels and the ground-truth surface. By leveraging the texel-level correspondences, our approach (Ours wo SR) achieves a substantial improvement in surface reconstruction accuracy, along with further gains in, both, accuracy and visual quality for novel view and pose synthesis.
Texel Super Resolution. Through surface-, vertex-, and texel-level alignment, our method achieves unprecedented quality in both rendering and geometry. However, for challenging cases, details like fine texture patterns and yarn structures cannot be adequately captured with Gaussian textures with the original resolution as illustrated in Fig. 10. Therefore, as mentioned in Sec. 3.6, we apply the Gaussian Texture super-resolution module to the Gaussian Textures, denoted as Ours. As can be seen from Fig. 10, by applying the Gaussian Texture super-resolution module, the tiny glyphs on the clothing, could be properly reconstructed. Moreover, it also leads to quantitative improvements on both novel view and novel pose synthesis task. It is worth mentioning that, since the network only models the residual of the Gaussian Splatting parameters w.r.t. the original Gaussian Textures, and employs a lightweight architecture, its computational overhead remains minimal (Ours wo SR runs at $21$ fps, while Ours runs at $18$ fps).
Figure 10. Qualitative ablation for texel super resolution. By applying texel super-resolution, our model could further preserve the tiny patterns on the clothing. Please zoom-in to better observe the details.
5. Applications
5.1. VR Telepresence
We implemented a VR demo to visualize the results generated by UMA . The system is built with Unity3D and integrates the gsplat Gaussian Splat (Ye et al., 2025) renderer. For each frame, the Gaussian Splatting results are precomputed and stored on disk. The Unity3D backend receives the VR headset’s pose in real time and renders the corresponding view, which is then streamed from the workstation to the headset. As shown in Fig. 11, users can walk around to inspect the virtual character.
Figure 11. VR Application. Users may inspect the dynamic avatar in VR headsets. The high-quality avatar geometry and appearance provides unprecedented immersive experiences, which may further boost the applications such as teleconference, remote assistance and so on.
5.2. Motion Editing
The animatable avatar generated by UMA paves the way for creating photorealistic, high-quality content, enabling users to animate the character with desired motions and render it from arbitrary viewpoints and zoom-in levels. To facilitate animation creation, we present UMA-Viewer, a tool that enables users to inspect recorded sequences as well as edit and visualize renderings of newly generated motion sequences. The front-end system is implemented with Viser (Yi et al., 2025) and runs on a personal computer. The backend is implemented in PyTorch, generating renderings on the fly based on the skeleton motion and camera parameters received from the front-end, and streaming results back at an interactive frame rate.
Inspecting recorded sequences. UMA-Viewer enables users to inspect characters driven by recorded motion sequences. Specifically, users can playback entire sequences from various camera viewpoints. For any given frame, the character can be freely rotated and zoomed in for detailed inspection. Furthermore, the system supports playback along user-defined camera trajectories, with adjustable zoom-in levels along the camera path.
Figure 12. User interface for UMA-Viewer. UMA-Viewer offers is an interactive system to inspect the recorded sequence and create novel motion sequences. (A) The control panel for changing the rendering settings. (B) The DOF panel editing and generating motion sequences.
Creating Novel Sequences. Beyond inspecting recorded motion sequences, UMA-Viewer also enables users to interactively create and examine novel motion sequences. Specifically, users can assign values to each degree of freedom (DOF) for newly created motion frames. Once the DOF is modified, the system will immediately visualize the character rendering.
Editing motion sequences frame by frame is a laborious and time-consuming task. To address this, UMA-Viewer allows users to efficiently generate motion sequences through key-framing. Specifically, for each DOF to be edited, users can specify the step size and the sequence length for motion sequence. As illustrated in Fig. 13, the resulting animatable character under the edited motion sequence exhibits plausible wrinkle dynamics and detailed textures.
Figure 13. Motion Editing Results. The rendering with motion generated with UMA-Viewer, note the clothing wrinkles caused by bending the elbow and raising the arm.
5.3. Motion Retargeting.
UMA is trained on high-quality, long multi-view sequences encompassing a diverse range of motions, enabling it to generalize effectively to novel poses. In Fig. 14, we present motion re-targeting results where all characters are driven by the same skeletal motion. Despite the motions being unseen during training, our method successfully renders photorealistic avatars with intricate appearance details and realistic wrinkles.
Figure 14. Motion retargeting. UMA enables animating different characters using the same skeletal motion depicted in the leftmost column, while faithfully preserving fine texture details and producing photorealistic wrinkles.
6. Limitations and Future Work
Though UMA makes clear improvements towards ultra-detailed human avatar modeling, there are still several open challenges left to solve. Since the clothed human avatar is driven with a single-layered person specific template, it does not support changing of the outfits. A layered representation which separates the modeling of garments and body could support the virtual try-on applications. Moreover, since the surface deformation of the garments is only conditioned on the skeletal motions, UMA currently does not support physical effects due to the external forces such as the interaction between clothing and objects. Besides, although UMA reconstructs high-fidelity surface geometry and apperance, which is crucial for applications for relighting. Later efforts could integrate the estimation of lights and materials to support the relighting of the ultra-high resolution rendering of clothed humans. Lastly, UMA primarily allows the user to control the skeletal motion, i.e., the body pose and hand poses. Future work could explore more expressive control including for example facial expressions.
7. Conclusion